# thanks: NJ and AC contributed equally to this project

### v1

@@ IV Nodal Line and Dirac Metal State of YMG
To identify the origin of the nodal lines, we consider the minimal set of four bands needed for a Dirac crossing. The generic low energy Hamiltonian describing such four bands is given by Tang et al. (2016); Fu and Kane (2007),
@@
Here, the five Dirac matrices, $\Gamma_{1}=\tau_{x}\otimes\sigma_{0}$, $\Gamma_{2}=\tau_{z}\otimes\sigma_{0}$, $\Gamma_{3}=\tau_{y}\otimes\sigma_{x}$, $\Gamma_{4}=\tau_{y}\otimes\sigma_{y}$, $\Gamma_{5}=\tau_{y}\otimes\sigma_{z}$ are constructed from the Pauli matrices $\tau_{\alpha}$ and $\sigma_{\alpha}$ with $\alpha=\{x,y,z\}$. The Pauli matrices represent the orbital ($\tau_{\alpha}$) and the spin ($\sigma_{\alpha}$) degrees of freedom. The coefficients $d_{i}(\textbf{k})$ (with $i=\{0,\cdots,5\}$) are real functions. All four bands become degenerate at a k point if $d_{i}(\textbf{k})=0$ for all $i$. Below, we analyze the impact of crystalline symmetry operations of magnetic YMG on the nodal line and band crossing points captured by this model.
@@
In presence of SOC with magnetization anisotropy along the $a$ axis Qiao et al. (2021), the $\mathcal{M}_{x}$ symmetry remains intact, while the $\mathcal{M}_{y}$ symmetry is broken. The anisotropy of the spin orientation in the orbital space leads to different crystalline symmetries, compared to the no SOC scenario (see Sec. V of SI for details Note (1)). Furthermore, YMG contains an additional two-fold $\mathcal{C}_{2y}$ rotation symmetry with the spin anisotropy along $a$. The $\mathcal{C}_{2y}$ symmetry forces the $d_{4}(k_{x}=0,k_{y},k_{z}=0)$ coefficient of the Hamiltonian to be zero on the $k_{y}$ axis (see Sec. V of the SI for details Note (1)). As discussed above, the $d_{1}$, and $d_{3}$ coefficients vanish on the $k_{x}=0$ plane owing to the ${\cal M}_{x}$ symmetry. Combining these restrictions, Dirac crossing can appear only when $d_{2}(k_{x}=0,k_{y},k_{z}=0)$, and $d_{5}(k_{x}=0,k_{y},k_{z}=0)$ is zero on the $k_{y}$ axis. Accordingly, we find two stable isolated Dirac crossings at $k_{y}=\pm 0.23$, equidistant from the $\Gamma$ point on the $k_{y}$ axis [see Fig. 4e]. Both Dirac points are pinned at $E_{\mathrm{F}}$. We highlight that the inclusion of SOC alters the symmetries of YMG, which leads to a Dirac nodal loop to Dirac crossings transition Tang et al. (2016); Singh et al. (2018a, b).
@@
In Fig. 4f, we show that the four bands forming the Dirac crossing carry different sets of eigenvalues for the $\mathcal{M}_{x}$ and $\mathcal{C}_{2y}$ symmetry operators. This stabilizes the Dirac crossings and protects them from level repulsion. Dirac Fermions are known to coexist with AFM ordering in AFe2As2 (A=Ba, Sr) Chen et al. (2017) and in CuMnAs Šmejkal et al. (2017), as well. In YMG, the two bands are gapped at all other $k$-points in the $k$-space as shown in the nodal plot in Fig. 4e. The $\mathcal{Z}_{2}$ number on the $k_{y}=0$ and $k_{y}=\pi/a$ planes are 0 and 1 respectively. That captures the non-trivial topological phase transition along the $k_{y}$ direction at the Dirac crossing in YMG, which is shown in the evolution of the wannier charge centre in Sec. V of the SI Note (1). The bulk topology leads to non-trivial surface states. We present the spin-projected spectral function of the Mn-terminated (001) surface along with the bulk states over $k_{x}-k_{y}$ plane in Fig. 4g. Owing to ${\cal P}{\cal T}$ symmetric Kramer’s degenerate bulk bands, the bulk spectral functions feature no spin-splitting. However, the (001) surface breaks the $\mathcal{PT}$ symmetry, and as a consequence, the surface states become spin-polarized. Remarkably, the spin polarization of YMG surface states is distinct from the spin-polarized surface states observed in other Dirac systems such as CuMnAs Tang et al. (2016), MTe2 (M=NiGhosh et al. (2019), PtYan et al. (2017); Politano et al. (2018), CoHu et al. (2023); Chakraborty et al. (2023)). More interestingly, the in-plane nearest neighbour surface Mn atoms are connected by $C_{4z}\mathcal{T}$ symmetry. Consequently, the spin-polarized surface features an even d-wave spin-polarized order in the momentum space (see Fig. 4g).
@@ V Conclusion
Despite their vast advantages in AFM spintronics and memory device applications, room-temperature metallic AFM candidates are scarce. We have demonstrated that YMG is an exfoliable layered antiferromagnetic Dirac metal at room temperature. Combining our ab-initio calculations with the semi-classical Monte Carlo approach, we show that the AFM metal state in YMG arises from the interplay of short-range correlation effects and the partial ($\pi,\pi$) nesting of the Fermi surface. Furthermore, the symmetry-protected topological Dirac crossings in bulk give rise to novel spin-polarized surface states with altermagnetic d-wave like spin texture, with no net magnetization.
Our findings motivate further investigations into transport phenomena, spintronic applications, and emergent physics at the intersection of magnetism, topology, and metallicity in YMG and related compounds. For example, the preserved ${\cal P}{\cal T}$ symmetry in bulk YMG suggests the potential realization of both dissipationless and dissipative intrinsic nonlinear Hall and longitudinal transport phenomenaMandal et al. (2024); Das et al. (2023), which could be leveraged for energy-efficient device applications. Further exploration of YMG and related van-der-Waals materials Hofmann et al. (2001); Chen et al. (2017); Malaman et al. (1994) could unlock new pathways for designing power-efficient, fast-switching spintronic devices.

### v2

@@ IV Nodal Line and Dirac Metal State of YMG
To identify the origin of the nodal lines, we consider the minimal set of four bands needed for a Dirac crossing. The generic low-energy Hamiltonian describing such four bands is given by Tang et al. (2016); Fu and Kane (2007),
@@
Here, the five Dirac matrices, $\Gamma_{1}=\tau_{x}\otimes\sigma_{0}$, $\Gamma_{2}=\tau_{z}\otimes\sigma_{0}$, $\Gamma_{3}=\tau_{y}\otimes\sigma_{x}$, $\Gamma_{4}=\tau_{y}\otimes\sigma_{y}$, $\Gamma_{5}=\tau_{y}\otimes\sigma_{z}$ are constructed from the Pauli matrices $\tau_{\alpha}$ and $\sigma_{\alpha}$ with $\alpha=\{x,y,z\}$. The Pauli matrices represent the orbital ($\tau_{\alpha}$) and the spin ($\sigma_{\alpha}$) degrees of freedom. The coefficients $d_{i}(\textbf{k})$ (with $i=\{0,\cdots,5\}$) are real functions. All four bands become degenerate at a k point if $d_{i}(\textbf{k})=0$ for all $i$. Below, we analyse the impact of crystalline symmetry operations of magnetic YMG on the nodal line and band crossing points captured by this model.
@@
In presence of SOC with magnetisation anisotropy along the $a$ axis Qiao et al. (2021), the $\mathcal{M}_{x}$ symmetry remains intact, while the $\mathcal{M}_{y}$ symmetry is broken. The anisotropy of the spin orientation in the orbital space leads to different crystalline symmetries, compared to the no SOC scenario (see Sec. V of SI for details Note (1)). Furthermore, YMG contains an additional two-fold $\mathcal{C}_{2y}$ rotation symmetry with the spin anisotropy along $a$. The $\mathcal{C}_{2y}$ symmetry forces the $d_{4}(k_{x}=0,k_{y},k_{z}=0)$ coefficient of the Hamiltonian to be zero on the $k_{y}$ axis (see Sec. V of the SI for details Note (1)). As discussed above, the $d_{1}$, and $d_{3}$ coefficients vanish on the $k_{x}=0$ plane owing to the ${\cal M}_{x}$ symmetry. Combining these restrictions, Dirac crossing can appear only when $d_{2}(k_{x}=0,k_{y},k_{z}=0)$, and $d_{5}(k_{x}=0,k_{y},k_{z}=0)$ is zero on the $k_{y}$ axis. Accordingly, we find two stable isolated Dirac crossings at $k_{y}=\pm 0.23$, equidistant from the $\Gamma$ point on the $k_{y}$ axis [see Fig. 4e]. Both Dirac points are pinned at $E_{\mathrm{F}}$. We highlight that the inclusion of SOC alters the symmetries of YMG, which leads to a Dirac nodal loop to Dirac crossings transition Tang et al. (2016); Thakur et al. (2017); Singh et al. (2018a, b).
@@
In Fig. 4f, we show that the four bands forming the Dirac crossing carry different sets of eigenvalues for the $\mathcal{M}_{x}$ and $\mathcal{C}_{2y}$ symmetry operators. This stabilises the Dirac crossings and protects them from level repulsion. Dirac Fermions are known to coexist with AFM ordering in AFe2As2 (A=Ba, Sr) Chen et al. (2017) and in CuMnAs Šmejkal et al. (2017), as well. In YMG, the two bands are gapped at all other $k$-points in the $k$-space as shown in the nodal plot in Fig. 4e. The $\mathcal{Z}_{2}$ number on the $k_{y}=0$ and $k_{y}=\pi/a$ planes are 0 and 1 respectively. That captures the non-trivial topological phase transition along the $k_{y}$ direction at the Dirac crossing in YMG, which is shown in the evolution of the wannier charge centre in Sec. V of the SI Note (1). The bulk topology leads to non-trivial surface states. We present the spin-projected spectral function of the Mn-terminated (001) surface along with the bulk states over $k_{x}-k_{y}$ plane in Fig. 4g. Owing to ${\cal P}{\cal T}$ symmetric Kramer’s degenerate bulk bands, the bulk spectral functions feature no spin-splitting. However, the (001) surface breaks the $\mathcal{PT}$ symmetry, and as a consequence, the surface states become spin-polarised. Remarkably, the spin polarisation of YMG surface states is distinct from the spin-polarised surface states observed in other Dirac systems such as CuMnAs Tang et al. (2016), MTe2 (M=NiGhosh et al. (2019), PtYan et al. (2017); Politano et al. (2018), CoHu et al. (2023); Chakraborty et al. (2023)). More interestingly, the in-plane nearest neighbour surface Mn atoms are connected by $C_{4z}\mathcal{T}$ symmetry. Consequently, the spin-polarised surface features an even d-wave spin-polarised order in the momentum space (see Fig. 4g).
@@ V Conclusion
Despite their vast advantages in AFM spintronics and memory device applications, room-temperature metallic AFM candidates are scarce. We have demonstrated that YMG is an exfoliable layered antiferromagnetic Dirac metal at room temperature. Combining our ab-initio calculations with the semi-classical Monte Carlo approach, we show that the AFM metal state in YMG arises from the interplay of short-range correlation effects and the partial ($\pi,\pi$) nesting of the Fermi surface. Furthermore, the symmetry-protected topological Dirac crossings in bulk give rise to novel spin-polarised surface states with altermagnetic d-wave like spin texture, with no net magnetisation.
Our findings motivate further investigations into transport phenomena, spintronic applications, and emergent physics at the intersection of magnetism, topology, and metallicity in YMG and related compounds. For example, the preserved ${\cal P}{\cal T}$ symmetry in bulk YMG suggests the potential realization of both dissipationless and dissipative intrinsic nonlinear Hall and longitudinal transport phenomenaChakraborty et al. (2022); Mandal et al. (2024); Das et al. (2023), which could be leveraged for energy-efficient device applications. Further exploration of YMG and related van-der-Waals materials Hofmann et al. (2001); Chen et al. (2017); Malaman et al. (1994) could unlock new pathways for designing power-efficient, fast-switching spintronic devices.